A cup of coffee was brewed in a container that was sealed. Three gaseous substances were obtained above the liquid. The gases were collected and separated into individual containers with the same temeprature (T), pressure (P), and volume (V). The first compound was weighed and gave value of 0.035 g. It was also analyzed and identified as H2O. The other compounds werefound to be N2 and O2. What is the weight of N2 and O2?
Answer #1

from ideal gas equation

PV= (w/M)RT

PV/RT = W/M

w = wt of water = 0.035 g

M = molarmass of water = 18 g/mol

PV/RT   = 0.035/18

   PV/RT = 0.00194

mass of N2(w) =

molarmass of N2 = 28 g/mol

0.00194 = w/28

w = 0.0543 g


mass of O2(w) =

molarmass of o2 = 32 g/mol

0.00194 = w/32

w = 0.06208 g